OpenStudy (anonymous):

how do you solve solve x^4+5x^2-6=0

OpenStudy (austinl):

\( x^4+5x^2-6=0\) I would factor this one out.

OpenStudy (anonymous):

plug x^2=t and solve for t and finally x

OpenStudy (debbieg):

Just remember that \(\Large x^4=(x^2)^2\) You could make an explicit substitution like \(\Large u=x^2\) to solve for u, then find x from that. But really you can probably do it without the substituion as the factoring is pretty straight-forward.

OpenStudy (anonymous):

The attached plot may help.

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."